# Fast and Robust Simulation-Based Inference With Optimization Monte Carlo

### v1

@@ Abstract
Bayesian parameter inference for complex stochastic simulators is challenging due to intractable likelihood functions. Existing simulation-based inference methods often require large number of simulations and become costly to use in high-dimensional parameter spaces or in problems with partially uninformative outputs. We propose a new method for differentiable simulators that delivers accurate posterior inference with substantially reduced runtimes. Building on the Optimization Monte Carlo framework, our approach reformulates stochastic simulation as deterministic optimization problems. Gradient-based methods are then applied to efficiently navigate toward high-density posterior regions and avoid wasteful simulations in low-probability areas. A JAX-based implementation further enhances the performance through vectorization of key method components. Extensive experiments, including high-dimensional parameter spaces, uninformative outputs, multiple observations and multimodal posteriors show that our method consistently matches, and often exceeds, the accuracy of state-of-the-art approaches, while reducing the runtime by a substantial margin.
@@ 1 Introduction
A variety of methods, collectively referred to as Likelihood-Free (LFI) or Simulation-Based (SBI) Inference (lintusaari2017fundamentals; sisson2018handbook; cranmer2020frontier; deistler2025simulation), have been developed to infer simulator parameters. These include sampling-based techniques, like Approximate Bayesian Computation (ABC, Marin2012), and approaches based on summary statistics (Chen2021a; Chen2023), parametric modeling (wood2010statistical; price2018bayesian), and ratio estimation (thomas2022likelihood; durkan2020contrastive). More recently, substantial progress has been driven by neural methods, which use deep networks to estimate the posterior (greenberg2019automatic; papamakarios2016fast; radev2020bayesflow; wildberger2023flow) or the likelihood (papamakarios2019sequential) function.
@@
Code is available at github.com/givasile/lfi.
@@ 2 Background and Motivation
Neural-based SBI methods are data-intensive, with their accuracy depending on access to large simulated datasets. This incurs computational costs on two fronts: generating massive datasets from complex simulators and training deep neural estimators on such volumes. Prior work (lueckmann2021benchmarking) has shown that it is the latter that often dominates runtime. Moreover, modern frameworks, like JAX, are used in SBI (dirmeier2024simulation) to accelerate simulation through vectorization, which can massively reduce the first cost but not the second. Appendix A expands on both points. This motivates us to develop a method that avoids the cost of training an external neural estimator.
@@ 2.1 ROMC framework
ROMC, introduced by ikonomov2020robust, improves upon Optimization Monte Carlo meeds2015optimization by addressing its robustness issues. It belongs to the larger family of ABC methods that approximate the likelihood function as the probability of simulating an output $\epsilon$-close to the observation,
@@ 3.2 Multiple iid observations
The rationale for choosing $q(\bm{\theta})$ in (12) as proposal distribution is as follows: it is a mixture of uniform distributions, so sampling and evaluating is easy. It seamlessly combines the contribution from multiple data points, allowing future data points to be included without re-computation. If each $q_{i}^{n}$ encloses the corresponding acceptance region $C_{\epsilon}^{i,n}$, then $q(\bm{\theta})>0$ in areas of the parameter space where $\prod_{n}^{N}\sum_{i}^{S}\mathbb{1}_{C_{\epsilon}^{i,n}}(\bm{\theta})>0$ and no posterior mass is discarded in the weighting in (14). Finally, regions of high likelihood are characterized by regions where multiple $C_{\epsilon}^{i,n}$ overlap. As the proposal distribution $q(\bm{\theta})$ consists of a mixture of the $q_{i}^{n}(\bm{\theta})$ with equal weights, areas where multiple $C_{\epsilon}^{i,n}$ overlap will be sampled more often, so that areas of high likelihood will be represented with more samples.
Figure 2, steps (ii-iii), illustrates the proposed procedure to obtain posterior samples. We have two observed data points and the $\bm{\theta}_{p}$ come from regions that match at least one observation; the blue-dotted area for $\mathbf{y}^{o}_{1}$ and the green-dotted region for $\mathbf{y}^{o}_{2}$. However, only the ones that match both will get a positive weight (red dots) and become posterior samples.
@@ 3.4 Limitations and Discussion
In the multiple-observations setting, R2OMC optimizes each observation independently. As a result, each proposed samples is $\epsilon$-close to at least one observation but not equally-close to others, and this cannot be improved since optimization is per observation. In such cases, accepting enough proposal samples ($w_{p}>0$ in Eq. (14)) requires increasing the $\epsilon$-threshold, which can broaden the posterior beyond the ground truth. However, under the assumption that all observations originate from a single (unknown) parameter configuration, many proposal samples are expected to render all observations plausible, so $\epsilon$ need not become large.
@@ 4 Experiments
We assess R2OMC on: (a) a Mixture of Gaussians (MoG) benchmark, (b) the synthetic SBI benchmark (lueckmann2021benchmarking) and (c) two image-based inference tasks. Together, they span diverse challenges, including high-dimensionality, distractors, multiple observations, and complex posteriors.
@@
Figure 4: SLCP pairwise posteriors for multiple observations. The four left panels show proposal samples per observation within $\epsilon$-distance, while the rightmost panel shows the subset of samples selected as closest across all observations using the R2OMC weighting scheme (Eq. 14).
@@ 4.2 SBI benchmark
Figure 5 reports the C2ST score versus runtime for R2OMC and the methods of lueckmann2021benchmarking on both SLCP variants. R2OMC approximates the posterior accurately (C2ST score 0.70.70.80.8) in some seconds, whereas other methods require substantially longer runtimes (often hours), because they require budgets of at least $10^{4}$ (often $10^{5}$) samples for similar performance. The gap is especially pronounced in the presence of distractors, where most methods fail to achieve C2ST scores below $0.8$ regardless of budget.
Figure 4 illustrates how R2OMC handles multiple observations. Proposal samples are first selected per observation (within $\epsilon$-distance) and then weighted according to Eq. 14 to retain those closest with respect to all observations. We speculate that this is why C2ST scores never reach the optimum of 0.5 and starts to flatten out: increasing the budget ensures samples are $\epsilon$-close to individual observations, but not necessarily to the joint set. In our case, all proposed samples satisfy $\epsilon<1$ per observation, yet the maximum distance across all observations reaches $\epsilon=5$.
@@
Figure 6 compares R2OMC to the methods of lueckmann2021benchmarking. R2OMC achieves a near-optimal C2STscore ($\approx 0.5$) in a few-seconds runtime, whereas competing methods require minutes to hours (budgets of $10^{5}$ samples) to reach comparable performance. The pairwise posterior further demonstrates that R2OMC accurately captures both crescent-shaped modes. Additional results versus budget are reported in Appendix D.
@@ 5 Conclusion
We have presented R2OMC, a new Bayesian inference method for differentiable simulator models that addresses two major challenges in simulation-based (likelihood-free) inference: scalability to high-dimensional parameter spaces and robustness to uninformative dimensions (distractors), all within fast runtimes.
Building on the ROMC framework, R2OMC recasts inference as a set of deterministic optimization problems solvable via gradient descent, enabling efficient inference in high-dimensional spaces. A gradient-based filtering step further masks distractors, while a novel adaptive importance sampling scheme selects posterior samples consistent with all available i.i.d. observations.
Our JAX-based implementation leverages auto-differentiation and vectorization to accelerate simulator evaluations, enabling many parameter–seed combinations to be executed at the cost of a single call. Across both novel experimental settings and the comprehensive SBI benchmark, R2OMC consistently achieves high accuracy—often approaching the optimal C2ST score of 0.5—while requiring only a fraction of the computational time of state-of-the-art methods.
@@ D.1.3 Results
For a more detailed view, Figures 9 ($\mathcal{S}_{\mathtt{base}}$), 10 ($\mathcal{S}_{\mathtt{base}}^{\mathtt{dist}}$), 11 ($\mathcal{S}_{\mathtt{MoG}}$), and 12 ($\mathcal{S}_{\mathtt{MoG}}^{\mathtt{dist}}$) report the mean C2ST score for each method across all dimensionalities and budgets.

### v2

@@ Abstract
Bayesian parameter inference for complex stochastic simulators is challenging due to intractable likelihood functions. Existing simulation-based inference methods often require large number of simulations and become costly to use in high-dimensional parameter spaces or in problems with partially uninformative outputs. We propose a new method for differentiable simulators that delivers accurate posterior inference with substantially reduced runtimes. Building on the Optimization Monte Carlo framework, our approach reformulates inference for stochastic simulators in terms of deterministic optimization problems. Gradient-based methods are then applied to efficiently navigate toward high-density posterior regions and avoid wasteful simulations in low-probability areas. A JAX-based implementation further enhances the performance through vectorization of key method components. Extensive experiments, including high-dimensional parameter spaces, uninformative outputs, multiple observations and multimodal posteriors show that our method consistently matches, and often exceeds, the accuracy of state-of-the-art approaches, while reducing the runtime by a substantial margin.
@@ 1 Introduction
A variety of methods, collectively referred to as Likelihood-Free (LFI) or Simulation-Based (SBI) Inference methods (lintusaari2017fundamentals; sisson2018handbook; cranmer2020frontier; deistler2025simulation; arruda2025diffusion), have been developed to infer simulator parameters. These include Approximate Bayesian Computation (ABC, Marin2012), summary statistic-based approaches (Chen2021a; Chen2023), parametric modeling (wood2010statistical; price2018bayesian; pacchiardi2024generalized), ratio estimation (thomas2022likelihood; durkan2020contrastive), and surrogate model optimisation (gutmann2016bayesian; jarvenpaa2025surrogate). Recently, neural-based methods have become prominent, with deep neural networks used to estimate the posterior (greenberg2019automatic; papamakarios2016fast; radev2020bayesflow; wildberger2023flow) or the likelihood (papamakarios2019sequential) function. State-of-the-art methods include Simformer (gloeckler2024allinone) and tabular-foundation-model SBI (vetter2025effortless), which achieve high accuracy by learning arbitrary conditionals of the joint distribution.
@@
Code for using R2OMC is publicly available at github.com/givasile/lfi. To reproduce the experiments exactly, use the branch github.com/givasile/lfi @ aistats2026.
@@ 2 Background and Motivation
Neural-based SBI methods are data-intensive, with their accuracy depending on access to large simulated datasets. This incurs computational costs on two fronts: generating massive datasets from complex simulators and training deep neural estimators on such volumes. Prior work (lueckmann2021benchmarking) has shown that it is the second cost, neural network training, that often dominates SBI. Moreover, modern frameworks, like JAX, are used in SBI (dirmeier2024simulation; quera2023blackbirds) to accelerate simulation through vectorization, which can massively reduce the first cost but not the second. Appendix A expands on both points. This motivates us to develop a method that avoids the cost of training an external neural estimator.
@@ 2.1 ROMC framework
ROMC, introduced by ikonomov2020robust, improves upon Optimization Monte Carlo of meeds2015optimization by addressing its robustness issues. It belongs to the larger family of ABC methods that approximate the likelihood function as the probability of simulating an output $\epsilon$-close to the observation,
@@ 3.2 Multiple iid observations
The rationale for choosing $q(\bm{\theta})$ in (12) as proposal distribution is as follows: it is a mixture of uniform distributions, so sampling and evaluating is easy. It seamlessly combines the contribution from multiple data points, allowing future data points to be included without re-computation. If each $q_{i}^{n}$ encloses the corresponding acceptance region $C_{\epsilon}^{i,n}$, then $q(\bm{\theta})>0$ in areas of the parameter space where $\prod_{n}^{N}\sum_{i}^{S}\mathbb{1}_{C_{\epsilon}^{i,n}}(\bm{\theta})>0$ and no posterior mass is discarded. Finally, regions of high likelihood are characterized by regions where multiple $C_{\epsilon}^{i,n}$ overlap. As the proposal distribution $q(\bm{\theta})$ consists of a mixture of the $q_{i}^{n}(\bm{\theta})$ with equal weights, areas where multiple $C_{\epsilon}^{i,n}$ overlap will be sampled more often, so that areas of high likelihood will be represented with more samples.
Figure 2, steps (ii–iii), illustrates the proposed procedure to obtain posterior samples $\bm{\theta}_{p}$. We have two observed data points and the $\bm{\theta}_{p}$ come from regions that match at least one observation; the blue-dotted area for $\mathbf{y}^{o}_{1}$ and the green-dotted region for $\mathbf{y}^{o}_{2}$. However, only the ones that match both will get a positive weight (red dots) and become posterior samples.
@@ 3.4 Limitations and Discussion
In the multiple-observations setting, R2OMC optimizes each observation independently. As a result, each proposed sample is $\epsilon$-close to at least one observation but not equally-close to others, and this cannot be improved since optimization is per observation. In such cases, accepting enough proposal samples ($w_{p}>0$ in Eq. (14)) requires increasing the $\epsilon$-threshold, which can broaden the posterior beyond the ground truth. However, under the assumption that all observations originate from a single (unknown) parameter configuration, many proposal samples are expected to make all observations plausible, so $\epsilon$ need not become large.
@@ 4 Experiments
We assess R2OMC on: (a) a Mixture of Gaussians (MoG) benchmark, (b) the synthetic SBI benchmark (lueckmann2021benchmarking), (c) two image-based inference tasks and (d) the Lotka-Volterra model. Together, they span diverse challenges, including high-dimensionality, distractors, multiple observations, complex posteriors and real-world applications.
@@
Figure 4: SLCP pairwise posteriors for multiple observations; blue: R2OMC samples, red: ground-truth samples. The four left panels show proposal samples per observation within $\epsilon$-distance, while the rightmost panel shows the subset of samples selected as closest across all observations using the R2OMC weighting scheme (Eq. 14).
@@ 4.2 SBI benchmark
Figure 5 reports the C2ST score versus runtime for R2OMC and the methods of lueckmann2021benchmarking on both SLCP variants. R2OMC approximates the posterior accurately (C2ST score 0.70.70.80.8) in a few seconds, whereas other methods require substantially longer runtimes (often hours), because they require budgets of at least $10^{4}$ (often $10^{5}$) samples for similar performance. The gap is especially pronounced in the presence of distractors, where most methods fail to achieve C2ST scores below $0.8$ regardless of budget.
Figure 4 illustrates how R2OMC handles multiple observations. Proposal samples are first selected per observation (within $\epsilon$-distance) and then weighted according to Eq. 14 to retain those closest with respect to all observations. We speculate that this is why C2ST scores never reach the optimum of 0.5 and start to flatten out: increasing the budget ensures samples are $\epsilon$-close to individual observations, but not necessarily to the joint set. In our case, all proposed samples satisfy $\epsilon<1$ per observation, yet the maximum distance across all observations reaches $\epsilon=5$.
@@
Figure 6 compares R2OMC to the methods of lueckmann2021benchmarking. R2OMC achieves a near-optimal C2STscore ($\approx 0.5$) in a few seconds, whereas competing methods require minutes to hours (budgets of $10^{5}$ samples) to reach comparable performance. The pairwise posterior further demonstrates that R2OMC accurately captures both crescent-shaped modes. Additional results versus budget are reported in Appendix D.
@@ 4.3 Image-based inference
4.4 Lotka-Volterra
Table 1: Lotka-Volterra SBC results (150 repetitions, budget = 1,000).
Parameter SBC $\chi^{2}_{100}$ $p$-value Cov. 50% Cov. 90% Cov. 95%
$\alpha$ (prey birth) 97.79 0.544 0.500 0.900 0.953
$\beta$ (predation) 84.32 0.870 0.500 0.927 0.940
$\gamma$ (predator death) 81.63 0.910 0.527 0.907 0.960
$\delta$ (reproduction) 89.71 0.760 0.480 0.907 0.987
Figure 8: Lotka-Volterra SBC summary. Left (table): $\chi^{2}$ uniformity tests ($k=100$ degrees of freedom) and empirical coverage suggest well-calibrated posteriors. Right: coverage calibration curve—empirical vs. nominal level across 150 repetitions; all parameters closely follow the diagonal.
Figure 9: Lotka-Volterra pairwise posterior for a representative observation. The black star marks the true parameter $\bm{\theta}^{*}$, which falls well within the posterior mass for all four parameters.
The Lotka-Volterra predator-prey model is a standard benchmark in the SBI literature (lueckmann2021benchmarking). It provides a realistic test case where the data-generating process is governed by a system of ODEs.
Setup. $X$ and $Y$ denote the prey and predator populations, with parameters $\alpha,\beta,\gamma,\delta$ governing their dynamics:
$\frac{dX}{dt}=\alpha X-\beta XY,\quad\frac{dY}{dt}=-\gamma Y+\delta XY$
Starting from $(X(0),Y(0))=(30,1)$ and integrating over $T=20$ iterations, the simulator produces a 20-dimensional time-series observation $\mathbf{y}=(x_{1,1},x_{2,1},\ldots,x_{1,10},x_{2,10})$ corrupted by log-normal noise. Parameters follow independent log-normal priors: $\alpha,\gamma\sim\mathrm{LogNormal}(-0.125,0.5)$, $\beta,\delta\sim\mathrm{LogNormal}(-3,0.5)$. The simulator is implemented in JAX for compatibility with R2OMC’s gradient-based optimisation.
Evaluation via SBC. Since our JAX-based ODE solver differs slightly from the Julia implementation of lueckmann2021benchmarking, ground-truth posterior samples are not available for a direct C2ST comparison. We therefore assess calibration via Simulation-Based Calibration (SBC, hermans2021trust), running 150 independent repetitions with a budget of 1,000 simulator calls each. SBC draws $\bm{\theta}^{*}\sim p(\bm{\theta})$, simulates $\mathbf{y}^{*}\sim p(\mathbf{y}|\bm{\theta}^{*})$, then draws $K$ posterior samples and records the rank of $\bm{\theta}^{*}$ among them. A well-calibrated posterior yields uniformly distributed ranks, which we test with a $\chi^{2}$ statistic. We additionally report empirical coverage: the fraction of repetitions in which $\bm{\theta}^{*}$ falls inside the $\alpha$-level credible interval. A calibrated posterior achieves empirical coverage $\approx\alpha$ for all levels.
Results. Figure 8 (Table 1 on the left and coverage calibration curve on the right) report the SBC $\chi^{2}$ statistics and the empirical coverage. All four parameters were consistent with a uniform distribution under the $\chi^{2}$ test (p >>\,0.54) and achieve near-nominal coverage, confirming well-calibrated posteriors within a budget of 1,000 calls. Figure 9 shows the pairwise posterior, where the true parameter $\bm{\theta}^{*}$ (black star) falls well within the posterior mass for all parameters, further indicating an accurate inference.
@@ 5 Conclusion
We have presented R2OMC, a new Bayesian inference method for differentiable simulator models that addresses two major challenges in simulation-based inference: scalability to high-dimensional parameter spaces and robustness to uninformative dimensions (distractors), all within fast runtimes.
Building on the ROMC framework, R2OMC recasts inference as a set of deterministic optimization problems solvable via gradient descent, enabling efficient inference in high-dimensional spaces. A gradient-based filtering step further masks distractors, while a novel adaptive importance sampling scheme selects posterior samples consistent with all available iid observations.
Our JAX-based implementation leverages auto-differentiation and vectorization, enabling many parameter–seed combinations to be executed at the cost of a single call. Across diverse benchmarks, R2OMC consistently achieves high accuracy while requiring only a fraction of the computational time of state-of-the-art methods.
@@ Acknowledgments
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Not Applicable]
    2. (b)
      Complete proofs of all theoretical results. [Not Applicable]
    3. (c)
      Clear explanations of any assumptions. [Not Applicable]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Yes]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
@@ D.1.3 Results
For a more detailed view, Figures 11 ($\mathcal{S}_{\mathtt{base}}$), 12 ($\mathcal{S}_{\mathtt{base}}^{\mathtt{dist}}$), 13 ($\mathcal{S}_{\mathtt{MoG}}$), and 14 ($\mathcal{S}_{\mathtt{MoG}}^{\mathtt{dist}}$) report the mean C2ST score for each method across all dimensionalities and budgets.